2. (8 points) Let F- try - vy2. Evaluate the line integralF dl using the f ollowing paths: a) (2 points) Going from 0, 0 to r 2,y 1 along the straight line y b) (2 points) Going from 0, 0 to r 2,y-1 along the parabola y c) (2 points) Going from 0,y 0 to 2,y 1 by going first along a vertical line from0,y0 to x-0, y-l then along the horizontal line x , y-i to x-2, y-1, d) (2 points) Going from 0,y to 2,y 1 ing the path (z(t), y()) where () 2t3,y)
Answer #1

Given,

overrightarrow{F} = hat{x}xy-hat{y}y^{2} .................................(1)

and,

I = int overrightarrow{F}.overrightarrow{dl}

or,   I = int left ( hat{x}xy-hat{y}y^{2} ight ).left ( hat{x}dx+hat{y}dy+hat{z}dz ight )

or, I = int xydx-y^{2}dy ....................................(2)

(a) Given,

y= rac{1}{2}x

or,   2 dy =-dr

And the range of x is from x= 0 to x = 2 .

Putting all the values in equation (2), we get

2 0

or,   0 3-8

or, I = rac{3}{8} left [ rac{x^{3}}{3} ight ]_{0} ^{2}

or,   5

or, I = 1

(b)  Given,

  14

so,   2 dy =-xdr

and the range of x is from x = 0 to x = 2.

Putting all the values in equation (2), we get

  2 2 2 0

or, 0

or, 16 32 6 5

or,   I = left [ 1 - rac{1}{3} ight ]

or,   I =rac{2}{3}

(c)  In this case, the integration is along the two paths : first along a vertical line from y = 0 (x= 0) to y = 1(x = 0) (i.e. along y - axis) and then along a horizontal line from x= 0 (y = 1) to x = 2 (y = 1) (i.e. parallel to the x - axis).

Let's denote the first path (i.e. along the y - axis) by 'A' and the second path (i.e. parallel to x -axis) by 'B' .

Integral along the path A :

The equation of the line (path of integration) will be , x = 0.

So, dx = 0

So, the integration will be

I_{A} =int_{0}^{1} 0- y^{2}dy

or, ム=- y dy

or, 3 2

or, I_{A} = - rac{1}{3}

Integral along the path B :

The equation of the line(path of integration) will be : y = 1 .

So, dy = 0

So, the integration will be

I_{B} = int_{0}^{2}left [ xdx - 0 ight ]

or, I_{B} = int_{0}^{2}xdx

or, I_{B} = left [ rac{x^{2}}{2} ight ] _{0} ^{2}

or,  I_{B} = 2

Hence the final result will be

I = I_{A}+I_{B}

or, I = - rac{1}{3} + 2 = rac{5}{3}

(d)  Given,

  x= 2t^{3} and y =t^{2}

So,   dx = 6 t^{2} dt and dy = 2tdt

And, for x = 0 ; t = 0 and for x = 2 ; t = 1 this condition (or limits) is also satisfied for y = 0 and y =1 .

So, I = int_{0}^{1} left [ 2t^{3} imes t^{2} imes 6t^{2}dt - left ( t^{2} ight )^{2} imes 2t dt ight ]

or,   I = int_{0}^{1} left [ 12t^{7} dt - 2 t^{5} dt ight ].

or,   I = left [ 12 , imes rac{t^{8}}{8} - 2 imes rac{t^{6}}{6} ight ]_{0} ^{1}

or, I = left [ 12 , imes rac{1}{8} - 2 imes rac{1}{6} ight ]

or, I = left [ rac{3}{2} - rac{1}{3} ight ]

or, I = rac{7}{6}

For any doubt please comment and please give an up vote. Thank you.